Skip to main content

AAVSOnet status, 2018-01-08

1 post / 0 new
HQA
HQA's picture
AAVSOnet status, 2018-01-08

Happy New Year, everyone!  Here is the current status of the AAVSOnet telescopes.

BSM_Berry  fully operational.

BSM_Hamren  fully operational.

BSM_NH  fully operational.  New filters and a grating have been added to the filter wheel, so the current complement is B,V,R,I,SU,SG,SR,SI,ZS,GRH.

BSM_NM  System has had some dust issues, which show up dramatically on the flat images.  We will probably have to send this camera back to SBIG to get cleaned.  Dust seems to flatten reasonably well, and so we are continuing to operate while discussing the problem.

BSM_S  fully operational.  A new filter wheel and filters to match BSM_NH are being shipped to BSM_S later this month.

SRO.  There was a computer failure on October 17.  While the system was rebuilt fairly quickly, some important parameters did not get set, and the pipeline processing has been using old flats.  This was only recently noticed.  We've taken on the challenge to reprocess all nights since October 17, so you may see some repeat images show up in your queue.  You might try comparing results with the new images; if no change, then don't worry about updating any submissions.  If there are changes, you might consider replacing AID submissions.

OC61.  In general, this system is working, though we have had some recent issues with dust, focus and tracking.  These are slowly getting fixed.

TMO61.  Fully operational.  The system was down for a couple of months while diagnosing some dome issues, but they seem to have been resolved.

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484